OpenStudy (anonymous):

Can someone check my simplification of this expression please (give me a sec to load it)

OpenStudy (anonymous):

\[\frac{5}{x}-\frac{2x-1}{x^2}+\frac{x+7}{x^3}\]

OpenStudy (anonymous):

and i came up with \[\frac{5x^6+5x^4-2x}{x^5}\]

OpenStudy (anonymous):

NO \[(5x^4-2x^2+2x+7)/x^3\]

OpenStudy (amistre64):

\[\frac{5}{x}*\frac{x^2}{x^2}-\frac{2x-1}{x^2}*\frac{x}{x}+\frac{x+7}{x^3}\] \[\frac{5x^2}{x^3}-\frac{x(2x-1)}{x^3}+\frac{x+7}{x^3}\] \[\frac{5x^2-x(2x-1)+x+7}{x^3}\] \[\frac{5x^2-2x^2-x+x+7}{x^3}\] \[\frac{3x^2+7}{x^3}\]

OpenStudy (amistre64):

.... forgot to switch a negative :)

OpenStudy (amistre64):

+2x up top

OpenStudy (anonymous):

no there is an error: -x+x is +x+x

OpenStudy (anonymous):

I don't know why i wanted to use the x^5 there, thanks guys let me work this out again

OpenStudy (anonymous):

bye ;)